OpenStudy (anonymous):

can anyone help with one quick math question? Piper buys 6 bottles of shampoo that have 42 ounces in each bottle. She also buys 3 bottles of shampoo that have 16 ounces in each bottle. What is the total number of ounces of shampoo that Piper buys?

OpenStudy (anonymous):

hi @AloneS XD can you help me figure this out?

alones (alones):

:) i'll try

OpenStudy (anonymous):

k :D

OpenStudy (anonymous):

I will give you a medal for trying :)

OpenStudy (anonymous):

540 ounces!

OpenStudy (anonymous):

thanks emily :D heres a medal, but can you explain how you got it?

OpenStudy (anonymous):

hmm one sec

OpenStudy (anonymous):

6 x 42 + 3 x 16!

OpenStudy (bamboo888):

300 ounces because it's 6x42+3x16=300. 540 in incorrect.

OpenStudy (anonymous):

oooh I get it so its 252+48=300

OpenStudy (anonymous):

oh oops! sorry about that guys!

OpenStudy (anonymous):

@Bamboo888 just making sure, the answer is 300?

alones (alones):

i think @Bamboo888 is right with her answer :)

OpenStudy (anonymous):

welp thanks I got an 86% on my sisters test, she's sick so I have to do the work for her XD poor little sister...

OpenStudy (bamboo888):

I'm sure 3x42=252 3x16=48 252+48= 300

OpenStudy (anonymous):

thanks again @Bamboo888

OpenStudy (bamboo888):

You're welcome. :)

OpenStudy (anonymous):

and @AloneS

alones (alones):

what ?? :)

OpenStudy (anonymous):

Did you take my medal away?

OpenStudy (anonymous):

no I didnt

OpenStudy (anonymous):

why do i have 0? Im kinda new so

OpenStudy (anonymous):

im closing this question now XD